Circular No. 6234

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. E. Matthews, Joint Astronomy Centre, Hilo, and Herzberg Insti-

tute of Astrophysics, Ottawa; and D. Jewitt and M. C. Senay, University of Hawaii, communicate: “We have detected the 230-GHz CO 2-1 rotational line in C/1995 O1, using the 15-m James Clerk Maxwell Telescope on Mauna Kea on Sept. 5, 7, 19 and 20 UT. Preliminary measurements of the Sept. 19 data yield line area 0.041 ± 0.006 K km s^{−1} and peak temperature 0.08 K. The line center is blueshifted by 0.35 ± 0.05 km/s with respect to the instantaneous geocentric velocity, indicating preferential sunward ejection of gas from the nucleus. Similar blueshifts were observed on the other three nights. Assuming a kinetic temperature of 10 K (as deduced for CO in 29P/Schwassmann-Wachmann 1 at a similar heliocentric distance; cf. IAUC 5929), we infer a CO production rate of 700 kg/s at r = 6.7 AU. The CO is presumably responsible for the creation of the extended dust coma seen in C/1995 O1, and thus for the overall brightness of this comet at large heliocentric distances. We expect C/1995 O1 to continue as a CO-driven comet until water sublimation dominates the volatile production at ∼ 3.5 AU (summer 1996).”

Z. Sekanina, Jet Propulsion Laboratory, has used the above production rate to calculate a correspondence to a sublimation area of 5 km^{2} if near the sub-solar point. This in turn suggests, if only 1 percent of the nucleus were active, that the nucleus’ size need not be larger than 10-15 km.

COMET 73P/SCHWASSMANN-WACHMANN 3 Total visual magnitude estimates suggest that this comet is indeed in

outburst (cf. IAUC 6215, 6227): Sept. 17.09 UT, 8.3: (A. Hale, Cloudcroft, NM, 0.2-m reflector); 18.43, 8.3 (J. Kobayashi, Kumamoto, Japan, 0.41-m reflector; 4′ coma); 21.13, 8.3: (C. S. Morris, Pine Mtn. Club, CA, 0.26-m reflector; low altitude; 30′ tail in p.a. 110^{o}).

V705 CASSIOPEIAE A. Retter and E.M. Leibowitz, Wise Observatory, Tel Aviv University,

report that observations of V705 Cas (N Cas 1993) over ten nights in Aug. and Sept. with the CCD camera (+ I filter) yield a light curve that shows clear periodic oscillations with a peak-to-peak amplitude of 0.05 mag and period 0.2280 ± 0.0005 day.

1995 September 20 Daniel W. E. Green